OpenStudy (anonymous):

A chemist needs to mix an 18% acid solution with a 45% acid solution to obtain a 12 liter mixture consisting of 36% acid. How many liters of each of the solutions must be used?

OpenStudy (anonymous):

a = liters of first solution b = liters of second solution a+b=12 and 0.18a+0.45b=0.36*12 0.18a+0.45b=4.32 Let's solve for a in the first equation. a+b=12 a=12-b Substitute it into the second. 0.18(12-b)+0.45b=4.32 2.16-0.18b+0.45b=4.32 0.27b=2.16 b=8 Because a+b=12, a+8=12, and a=4. You need 4L of 18% and 8L of 45%.